OpenStudy (anonymous):

Evaluate if possible: Square root of the fraction 64/81?

OpenStudy (anonymous):

8/9

OpenStudy (saifoo.khan):

\[\sqrt{\frac{64}{81}}\]

OpenStudy (saifoo.khan):

that?

OpenStudy (anonymous):

even if it was \[\sqrt{64}/81\] it would still be 8/9

OpenStudy (saifoo.khan):

sungirl?

OpenStudy (anonymous):

Sry! I'm working on my laptop and for whatever reason can't reply on here with it... so I ran up to my other computer to not seem rude! Yes, that is the correct form, and the right answer as well. Thanks!!!

OpenStudy (saifoo.khan):

Noprobs.